### v1

@@ Abstract.
We study the $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and Sobolev regularity for appropriate weak formulations of the (stationary) Laplace–Beltrami, Bochner Laplace, tangent Stokes, and tangent Navier–Stokes problems on compact, connected $d$-dimensional manifolds without boundary of class $C^{m}$ ($m\geq 2$) embedded in $\mathrm{R}^{d+1}$, in terms of the regularity of the source terms and the manifold. Firstly, we localize and translate the Laplace–Beltrami problem into flat domains to establish $\mathrm{L}^{p}$-based ($2\leq p<\infty$) Sobolev regularity results as a consequence of the powerful Calderón-Zygmund theory. From then onwards, our strategy is purely “variational”, building upon the abstract functional analytic idea of “duality” and classical results such as the Banach–Nečas–Babuška theorem and the generalized Babuška–Brezzi theory in reflexive Banach spaces to analyze and relate to each other different strong, weak and ultra-weak formulations for the aforementioned problems. In particular, by exploiting the lack of manifold boundary, we are able to decouple the velocity and the pressure unknowns in the tangent Stokes problem to establish their higher-regularity $\mathbf{W}^{k,p}\times\mathrm{W}^{k-1,p}$ ($k\geq 2$) as a consequence of the $\mathrm{L}^{p}$-based well-posedness and regularity theory for the Laplace–Beltrami and Bochner–Laplace operators. We finish this work by applying this newly derived regularity theory to: i) define and appropriate Stokes operator and study its spectral and regularity properties, ii) establish $\mathrm{L}^{p}$-based ($1<p<\infty$) existence and well-posedness for the (stationary) tangent Oseen and Navier–Stokes equations via a Galerkin method and bootstrapping, and higher $\mathrm{L}^{p}$-based Sobolev regularity results. iii) transfer our regularity results to other choices of vector Laplace operators.
@@
PDEs on manifolds, Sobolev regularity, well-posedness, Navier–Stokes, Stokes operator, Leray projector, duality, Banach–Nečas–Babuška, Babuška–Brezzi, inf-sup, Fredholm alternative, Galerkin method, fixed-point theory
@@ 1. Introduction
Let $\Gamma$ be a $d$-dimensional ($d\geq 2$) compact and connected $C^{2}$ manifold without boundary embedded in $\mathrm{R}^{d+1}$ and denote by $\boldsymbol{\nu}=(\nu_{i})_{i=1}^{d+1}$ its outward unit normal. Throughout this work, we are concerned with the following four problems:
      Laplace–Beltrami $-\Delta_{\Gamma}u$ $=f,$ (LB)
@@
where $f,g$ are given scalar fields with vanishing mean on $\Gamma$ and $\mathbf{f}$ is an everywhere tangential to $\Gamma$ vector field. The operators $\nabla_{\Gamma}$ and $\operatorname{div}_{\Gamma}$ respectively denote the covariant gradient and covariant divergence, and $\Delta_{\Gamma}:=\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ is the Laplace–Beltrami operator; see Section 2.1. The operator $\Delta_{B}:=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ is usually referred to as the Bochner Laplacian [40, §3.2], [64] or rough Laplacian [37], where $\mathbf{P}:=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ is the projection operator onto the tangent plane to $\Gamma$. The unknowns in (1.1) are the scalar functions $u:\Gamma\rightarrow\mathrm{R}$ and $\uppi:\Gamma\rightarrow\mathrm{R}$ with vanishing mean and the tangential vector field $\mathbf{u}:\Gamma\rightarrow\mathrm{R}^{d+1}$.
The problems (1.1) are of great importance as prototypical constituents of the modeling of a great variety of physical phenomena on thin films. For instance, the Laplace–Beltrami operator is a useful tool for image and shape processing [78] and in the modeling of surfactants [71]. In turn, (S) and (NS) are the basis in the study of fluid flows, with paramount theoretical [25, 51, 50, 49, 52] and practical importance, e.g. [43, 55, 22]. The governing system of geometric partial differential equations may be derived either by a formal asymptotic expansion [26, 59, 60] or by a rigorous approach [49, 53].
This work is devoted to establishing $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and Sobolev regularity results for appropriate weak formulations of problems (1.1) in terms of the regularity of the data ($f$, $\mathbf{f}$ and $g$) and manifold $\Gamma$. We stress that, unlike what is typically done in the differential geometry literature, we do not assume $\Gamma$ to be necessarily of class $C^{\infty}$, but rather we specify in each statement the required regularity on $\Gamma$. As far as the authors are concerned, our work is the first one to provide a fully-fledged well-posedness and regularity theory for problems (1.1), valid for any integrability parameter $1<p<\infty$ and manifolds of limited regularity and arbitrary intrinsic dimension $d\geq 2$. Additionally, we consider our exposition to be accessible to people without strong foundations on Riemannian geometry, hence serving as a good future reference for works in applied fields.
@@
For $\Omega\subseteq\mathrm{R}^{d}$ a bounded Lipschitz domain, we adopt the standard notation for Lebesgue spaces $\mathrm{L}^{q}(\Omega)$ and Sobolev spaces $\mathrm{W}^{m,q}(\Omega)$ and $\mathrm{W}^{m,q}_{0}(\Omega)$ with $m\geq 0$ and $q\in[1,\infty)$, whose corresponding norms are denoted respectively by $\|\cdot\|_{0,q;\Omega}$ and $\|\cdot\|_{m,q;\Omega}$. We also write $\mathrm{W}^{0,q}(\Omega)=\mathrm{L}^{q}(\Omega)$ and $\mathrm{H}^{m}(\Omega):=\mathrm{W}^{m,2}(\Omega)$ with corresponding norm $\|\cdot\|_{m,\Omega}:=\|\cdot\|_{m,2;\Omega}$ For any scalar function space $\mathrm{V}$ defined on $\Omega$, we denote by $\mathbf{V}$ and $\mathbb{V}$ its $d$-vectorial and $d\times d$ tensorial counterparts, and we use the same notation to denote their norms.
We will use a similar notation for the manifold Sobolev spaces $\mathrm{W}^{m,q}(\Gamma)$ defined on a $d$-dimensional ($d\geq 2$) compact and connected sufficiently regular manifold $\Gamma$ without boundary embedded in $\mathrm{R}^{d+1}$ (these spaces will be properly introduced in Section 2.2 below). For any scalar function space $\mathrm{V}$ defined on $\Gamma$, we denote by $\mathbf{V}$ and $\mathbb{V}$ its $(d+1)$-vectorial and $(d+1)\times(d+1)$ tensorial counterparts, and use the same notation for their respective norms. Moreover, we denote by $\mathrm{W}^{m,q}_{\#}(\Gamma)$ the closed subset of scalar-valued functions in $\mathrm{W}^{m,q}(\Gamma)$ that satisfy $\int_{\Gamma}v=0$, whereas $\mathbf{W}^{m,q}_{t}(\Gamma)$ denotes the space of vector-valued functions $\mathbf{v}:\Gamma\rightarrow\mathrm{R}^{d+1}$ that belong to $\mathbf{W}^{m,q}(\Gamma)$ and are a.e. tangential to $\Gamma$ (i.e. $\mathbf{v}\cdot\boldsymbol{\nu}=0$).
@@
Throughout this article, for any integrability parameter $p\in(1,\infty)$, we denote by $p^{*}:=\frac{p}{p-1}$ its conjugate exponent. Notice that $p^{**}=p$. Also, $1<p<2$ if and only if $p^{*}>2$.
@@ 1.1. Background
The involved differential operators in (1.1) satisfy integration-by-parts formulas that highly resemble their flat-domain counterparts [11, Ap. B], [24, §2] (see also [10, Lemma 1]). Said identities naturally induce a set of $\mathrm{L}^{2}$-based weak formulations that we now proceed to review.
Laplace–Beltrami problem
For the Laplace–Beltrami problem, its weak formulation reads: given $f\in(\mathrm{H}^{1}_{\#}(\Gamma))^{\prime}$, find $u\in\mathrm{H}^{1}_{\#}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}v=\langle f,v\rangle,\qquad\forall v\in\mathrm{H}^{1}_{\#}(\Gamma).$ (1.2)
For $\Gamma$ of class $C^{2}$ and of any intrinsic dimension $d\geq 2$, the well-posedness of (1.2) is a consequence of the Poincaré-Friedrichs inequality [10, Lemma 2] (see also [24, Theorem 2.12] for the result for $\Gamma$ of class $C^{3}$) and the Lax–Milgram theorem.
In flat bounded domains, a-priori estimates in general $\mathrm{W}^{m,p}$-norms for the scalar Laplace problem are a consequence of the Calderon–Zygmund theory applied to the integral representation of solutions via the Newtonian potential. For the case $m\geq 2$ we refer to [35, Chapter 9] and for the case $m=1$ we refer to [46, Section 4, Theorem 1] (see also [16, Theorem 1.5]). Said results transfer straightforwardly to the vector Laplace operator, as it corresponds to the componentwise application of the scalar Laplacian.
Regarding the higher Sobolev regularity of (LB) we refer to [10, Lemma 3] and [24, Theorem 3.3], which provides $\mathrm{H}^{2}$-regularity of the solution provided the right-hand side function $f$ lies in $\mathrm{L}^{2}(\Gamma)$ and $\Gamma$ is of class $C^{2}$. The proof follows from translating (1.2) into a general elliptic PDE whose $\mathrm{H}^{2}$-interior regularity is provided by classical results, e.g. [28, p. 327, Theorem 1].
In the context of differential forms, Hodge theory [54, §7.4] provides $\mathrm{W}^{2,p}$-regularity ($1<p<\infty$) for solutions of the Laplace–Beltrami problem—for 0-forms, i.e. scalar functions, the Hodge Laplacian coincides with the Laplace–Beltrami operator—in manifolds of class $C^{1,1}$.
@@
The well-posedness of (1.3) also follows from a Poincaré-type inequality for the covariant gradient $\nabla_{\Gamma}$, which has only been established for $\Gamma$ of class $C^{\infty}$ and of intrinsic dimension $d=2$ [37] (although, with a small gap in its proof). We fix this gap and extend the result to $C^{m}$, $m\geq 3$, manifolds of arbitrary dimension $d\geq 2$ in Section 3.3.
@@
We find it important to point out that different, but strongly related, second-order linear differential operators than the Bochner Laplacian $\Delta_{B}:=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ are considered throughout the vast differential geometry literature. For instance, in the context of differential forms, the Hodge Laplacian $\Delta_{H}=-(\,{\mathrm{d}}^{*}\,{\mathrm{d}}+\,{\mathrm{d}}\,{\mathrm{d}}^{*})$, defined in terms of the exterior derivative $\mathrm{d}$} and coddiferential $\mathrm{d}$$^{*}$, maps $1$-forms into $1$-forms; we refer to [72, p. 601] for its definition on general compact Riemannian manifolds. By means of the musical isomorphism, we also understand $\Delta_{H}$ as acting on tangent vector fields $\mathbf{v}:\Gamma\rightarrow\mathrm{R}^{d+1}$. Moreover, and primarily related to fluid flow problems, the so-called Surface Diffusion operator $\Delta_{S}:=\mathbf{P}\operatorname{div}_{\Gamma}2D_{\Gamma}$ (where $D_{\Gamma}:=\frac{1}{2}(\nabla_{\Gamma}+\nabla_{\Gamma}^{T})$ is the symmetrized covariant gradient) is considered to be the most physically d diffusion operator in the context of fluid flows [25, 42, 40]. We stress that none of these three operators coincide with the componentwise application of the scalar Laplace Beltrami operator $\Delta_{\Gamma}$.
@@
For any intrinsic dimension $d\geq 2$, the Bochner–Laplace, Hodge and Surface Diffusion operators are related via the identities [40, eq. (2.10)], [74, p. 583, eq. (3.56)]: for each $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$
$\Delta_{S}\mathbf{v}$ $=\Delta_{B}\mathbf{v}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v},$ (1.4)
$\Delta_{H}\mathbf{v}$ $=\Delta_{B}\mathbf{v}-(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}.$
@@
Here we have used that $\operatorname{Ric}(\mathbf{v})=(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}$, which follows from elementary manipulations of the Riemannian Curvature tensor (see (3.28) and (3.30) below), or alternatively, by comparing [73, p. 347, eq. (4.26)] with [40, eq. (2.10)].
@@
From (1.4) it is apparent that the Hodge Laplacian $\Delta_{H}$ and Bochner Laplacian $\Delta_{B}$ operators only differ on a zeroth-order term depending only on the curvatures of $\Gamma$. In turn, under the additional manifold incompressibility $\operatorname{div}_{\Gamma}=0$ assumption, then such relation also holds for the surface diffusion operator $\Delta_{S}$. For manifolds embedded in $\mathrm{R}^{3}$, the Hodge Laplacian can be rewritten as $\Delta_{H}=\operatorname{\boldsymbol{\operatorname{curl}}}_{\Gamma}\operatorname{curl}_{\Gamma}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}$. We refer to [40, §3.2] and [62, §4.1] and their references therein for a thorough discussion about these operators, including their historical origin and their relation with each other.
@@
The previously mentioned Hodge theory [54, §7.4] also provides $\mathrm{W}^{2,p}$-regularity ($1<p<\infty$) for solutions of the Hodge Laplacian $\Delta_{H}$ operators (for $1$-forms) in manifolds of class $C^{2,1}$. By virtue of the musical isomorphism, we expect the result on $\Delta_{H}$ to extend to tangent vector fields, although we are unable to pinpoint a specific reference for the proof. Consequently, said regularity would also be enjoyed by the Bochner Laplacian $\Delta_{B}$, as it only differs from the Hodge Laplace $\Delta_{H}$ in lower-order terms (cf. (1.4)).
@@
If we consider the Hodge Laplacian $\Delta_{H}$ or the Surface Diffusion $\Delta_{S}$ in lieu of the Bochner Laplacian $\Delta_{B}$, different Stokes systems arise; see for instance, [40, §§4-5] for $\Delta_{S}$, and [72, §17.A] for $\Delta_{H}$.
The weak formulation for the surface diffusion operator $\Delta_{S}$ [40, eq. (4.5)] replaces the diffusion term $\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$ in (1.5) by $\int_{\Gamma}D_{\Gamma}(\mathbf{u}):D_{\Gamma}(\mathbf{v})$. Consequently, the flow velocity $\mathbf{u}$ needs to be sought in the orthogonal complement of the non-trivial kernel of the operator $D_{\Gamma}$ so as to ensure uniqueness. The well-posedness of said formulation follows from a straightforward application of the Babuška–Brezzi theory (see, e.g. [30, Th. 2.1]), which hinges on a Korn-type inequality [40, Lemma 4.1] and an inf-sup condition on the covariant divergence $\operatorname{div}_{\Gamma}$ [40, Lemma 4.2]. As expected, the well-posedness of (1.5) will depend on the same Poincaré inequality (instead of a Korn inequality) for the covariant gradient $\nabla_{\Gamma}$ in the space $\mathbf{H}^{1}_{t}(\Gamma)$ used to establish the well-posedness of (1.3).
In flat bounded domains, estimates in $\mathrm{W}^{m,p}$-norms for Stokes systems, provided such regular-enough solutions exist, are a consequence of the results by Agmon, Douglis and Nirenberg [23, 2], whose proofs are built upon techniques of integral representation and potential theory. The existence of such solutions was proved in [17] for $d=3$, while the case $d=2$ is much simpler as it reduces to a biharmonic problem [75, Chapter 1, Proposition 2.3]. We refer to [75, Chapter 1, §2.5] for a thorough discussion on the regularity of the Stokes system and its early historical development.
The special case of $\mathrm{L}^{2}$-based (i.e. $p=2$) estimates allows alternative proofs that circumvent the use of potential theory; among them we find [32, 3, 6, 34]. Ghidaglia [32] (see also [19, §3]) argues based on difference quotients and flattening of the boundary, while Amrouche and Girault [3] develop a Helmholtz decomposition approach of the space $\mathbf{W}^{m,p}(\Omega)\cap\mathbf{W}^{1,p}_{0}(\Omega)$. Finally, we refer to Giaquinta and Modica [34] for a clear and self-contained discussion of the $\mathrm{L}^{2}$-based Sobolev and Hölder regularity of nonlinear Stokes systems with general coefficients.
We highlight the particularly elegant argument by Beirão da Veiga [6] (later extended by Berselli [9]) that, via the “method of continuity” (homotopy argument) [35, §9, Theorem 5.2], transfers the $\mathbf{W}^{2,p}(\Omega)$-well-posedness of the Poisson’s equation to $\mathbf{W}^{2,p}(\Omega)\times\mathrm{W}^{1,p}(\Omega)$.

### v2

@@ Abstract.
This paper is the second part of a two-paper series, initiated in [5] for scalar PDEs on hypersurfaces, and is concerned with the well-posedness and $\mathrm{L}^{p}$-based Sobolev regularity of vector-valued PDEs of interest in fluid dynamics. This family of PDEs includes the (stationary) Bochner Laplace, tangent Stokes and Oseen, and tangent Navier–Stokes equations. We present several strong, weak and ultra-weak formulations of these problems on compact, connected $d$-dimensional manifolds without boundary embedded in $\mathrm{R}^{d+1}$. We prove $\mathrm{W}^{m,p}$-regularity for any $p\in(1,\infty)$ for manifolds of minimal regularity $C^{m+1}$ or $C^{m,1}$ for $m\geq 1$. Building upon the $\mathrm{L}^{p}$-based scalar elliptic theory from [5], we develop a parametrization-free and purely variational approach that resorts to classical results such as the Banach–Nečas–Babuška theorem and the generalized Babuška–Brezzi theory in reflexive Banach spaces. In particular, by exploiting the manifold closedness, we decouple the velocity and pressure variables in the tangent Stokes problem to establish their higher-regularity $\mathbf{W}^{m,p}\times\mathrm{W}^{m-1,p}$ ($m\geq 2$) as a consequence of the $\mathrm{L}^{p}$-based well-posedness and regularity theory for the Laplace–Beltrami and Bochner–Laplace operators. We study spectral and regularity properties of an appropriate Stokes operator, and apply them to show existence of solutions for the Navier–Stokes equations for $p=2$ and $d\leq 4$. We next extend the well-posedness to $p>2$ and prove higher-order $\mathrm{L}^{p}$-based regularity. We finally examine alternative choices to the Bochner Laplace operator that are useful in fluid dynamics.
@@
PDEs on manifolds, Sobolev regularity, well-posedness, Navier–Stokes, Stokes operator, Leray projector, duality, Banach–Nečas–Babuška, Babuška–Brezzi, inf-sup, Fredholm alternative, Galerkin method, fixed-point
@@ 1. Introduction
Let $\Gamma$ be a $d$-dimensional ($d\geq 2$) compact, connected and sufficiently regular (to be specified later) manifold without boundary embedded in $\mathrm{R}^{d+1}$ and denote by $\boldsymbol{\nu}=(\nu_{i})_{i=1}^{d+1}$ its outward unit normal. Throughout this work, we are concerned with the following three vector-valued problems:
@@
where $g$ is a given scalar field with vanishing mean on $\Gamma$ and $\mathbf{f}$ is a vector field everywhere tangential to $\Gamma$. The operators $\nabla_{\Gamma}$ and $\operatorname{div}_{\Gamma}$ respectively denote the covariant gradient and covariant divergence; see Section 2.1 below and, e.g., [5, Sections 2-3]. The operator $\Delta_{B}:=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ is usually referred to as the Bochner Laplacian [36, §3.2], [56] or rough Laplacian [34], where $\mathbf{P}:=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ is the projection operator onto the tangent plane to $\Gamma$. The unknowns in (1.1) are the tangential vector field $\mathbf{u}:\Gamma\rightarrow\mathrm{R}^{d+1}$ and the scalar field $\uppi:\Gamma\rightarrow\mathrm{R}$ with vanishing mean.
The problems (1.1) are of great importance as prototypical constituents of the modeling of a great variety of physical phenomena on thin films. For instance, (S) and (NS) are the basis in the study of fluid flows, with paramount theoretical [24, 45, 44, 43, 46] and practical importance, e.g. [38, 49, 21]. The governing system of geometric partial differential equations may be derived either by a formal asymptotic expansion [25, 51, 52] or by a rigorous approach [43, 47].
This work is devoted to establishing $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and Sobolev regularity results for appropriate weak formulations of problems (1.1) in terms of the regularity of the data ($\mathbf{f}$ and $g$) and manifold $\Gamma$. We stress that, unlike what is typically done in the differential geometry literature, we do not assume $\Gamma$ to be necessarily of class $C^{\infty}$, but rather we specify in each statement the required regularity on $\Gamma$. As far as the authors are concerned, our work is the first one to provide a fully-fledged well-posedness and regularity theory for problems (1.1), valid for any integrability parameter $p\in(1,\infty)$ and manifolds of minimal regularity and arbitrary intrinsic dimension $d\geq 2$. Additionally, we consider our exposition to be accessible to people without strong foundations on Riemannian geometry, hence serving as a bridge to applied fields.
@@
We now introduce function spaces on the manifold $\Gamma$; these spaces are succintly discussed in Section 2.1 below (see [5, Section 3] for further details). We use the notation $\mathrm{L}^{q}(\Gamma)$ and $\mathrm{W}^{m,q}(\Gamma)$ for the scalar Lebesgue and Sobolev spaces with integrability $q\in[1,\infty)$ and differentiability $m\geq 0$; we denote the corresponding norms respectively by $\|\cdot\|_{0,q;\Gamma}$ and $\|\cdot\|_{m,q;\Gamma}$. We also write $\mathrm{W}^{0,q}(\Gamma)=\mathrm{L}^{q}(\Gamma)$ and $\mathrm{H}^{m}(\Gamma):=\mathrm{W}^{m,2}(\Gamma)$ with corresponding norm $\|\cdot\|_{m,\Gamma}:=\|\cdot\|_{m,2;\Gamma}$. For any scalar function space $\mathrm{V}$ defined on $\Gamma$, we denote by $\mathbf{V}$ and $\mathbb{V}$ its $(d+1)$-vectorial and $(d+1)\times(d+1)$ tensorial counterparts, and we use the same notation for their norms. Moreover, we denote by $\mathrm{W}^{m,q}_{\#}(\Gamma)$ the closed subset of scalar-valued functions in $\mathrm{W}^{m,q}(\Gamma)$ that satisfy $\int_{\Gamma}v=0$, whereas $\mathbf{W}^{m,q}_{t}(\Gamma)$ denotes the space of vector-valued functions $\mathbf{v}:\Gamma\rightarrow\mathrm{R}^{d+1}$ that belong to $\mathbf{W}^{m,q}(\Gamma)$ and are a.e. tangential to $\Gamma$ (i.e. $\mathbf{v}\cdot\boldsymbol{\nu}=0$).
@@
Throughout this article, for any integrability parameter $p\in(1,\infty)$, we denote by $p^{*}:=\frac{p}{p-1}$ its conjugate exponent. Notice that $p^{**}=p$. Also, $p\in(1,2)$ if and only if $p^{*}\in(2,\infty)$.
@@ 1.1. Background
The differential operators involved in (1.1) satisfy integration-by-parts formulas that highly resemble their flat-domain counterparts [11, Ap. B], [23, §2] (see also [10, Lemma 1]). Said identities naturally induce a set of $\mathrm{L}^{2}$-based weak formulations that we now review.
@@
The well-posedness of (1.2) follows from a Poincaré-type inequality for the covariant gradient $\nabla_{\Gamma}$, which has only been established for $\Gamma$ of class $C^{\infty}$ and of intrinsic dimension $d=2$ [34] (although, with a small gap in its proof). We fix this gap and extend the result to $\Gamma$ of class $C^{2}$ and arbitrary dimension $d\geq 2$ in Section 3.1.
@@
We find it important to point out that second-order linear differential operators other than the Bochner Laplacian $\Delta_{B}:=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ are considered throughout the vast differential geometry literature. For instance, in the context of differential forms, the Hodge Laplacian $\Delta_{H}=-(\,{\mathrm{d}}^{*}\,{\mathrm{d}}+\,{\mathrm{d}}\,{\mathrm{d}}^{*})$, defined in terms of the exterior derivative $\mathrm{d}$} and coddiferential $\mathrm{d}$$^{*}$, maps $1$-forms into $1$-forms; we refer to [60, p. 601] for its definition on general compact Riemannian manifolds. By means of the musical isomorphism, we also understand $\Delta_{H}$ as acting on tangent vector fields $\mathbf{v}:\Gamma\rightarrow\mathrm{R}^{d+1}$. Moreover, the so-called Surface Diffusion operator $\Delta_{S}:=\mathbf{P}\operatorname{div}_{\Gamma}2D_{\Gamma}$ (where $D_{\Gamma}:=\frac{1}{2}(\nabla_{\Gamma}+\nabla_{\Gamma}^{T})$ is the symmetrized covariant gradient) is considered to be the most physically relevant diffusion operator in the context of fluid flows [24, 37, 36]. We stress that none of these operators coincide with the componentwise application of the scalar Laplace Beltrami operator $\Delta_{\Gamma}$ studied in [5], whence the theory of (1.2) is not a simple consequence of [5].
@@
For any intrinsic dimension $d\geq 2$, the Bochner–Laplace $\Delta_{B}$, Hodge $\Delta_{H}$ and Surface Diffusion $\Delta_{S}$ operators are related via the identities [36, eq. (2.10)], [62, p. 583, eq. (3.56)]: for each $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$
$\Delta_{S}\mathbf{v}=\Delta_{B}\mathbf{v}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v},\qquad\Delta_{H}\mathbf{v}=\Delta_{B}\mathbf{v}-(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}.$ (1.3)
@@
Here we have used that the Ricci curvature tensor reads $\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2}$ in terms of the Weingarten map $\mathbf{B}:=\nabla_{\Gamma}\boldsymbol{\nu}$, which in turn encodes the principal curvatures of $\Gamma$. This follows from elementary manipulations of the Riemann Curvature tensor (see Appendix C below), or alternatively, by comparing [61, p. 347, eq. (4.26)] with [36, eq. (2.10)].
@@
From (1.3) it is apparent that the Hodge Laplacian $\Delta_{H}$ and Bochner Laplacian $\Delta_{B}$ operators only differ on a zeroth-order term $(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}$. Moreover, such a relation also holds for the surface diffusion operator $\Delta_{S}$ under the additional manifold incompressibility assumption $\operatorname{div}_{\Gamma}\mathbf{v}=0$. For manifolds embedded in $\mathrm{R}^{3}$, the Hodge Laplacian can be rewritten as $\Delta_{H}=\operatorname{\boldsymbol{\operatorname{curl}}}_{\Gamma}\operatorname{curl}_{\Gamma}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}$. We refer to [36, §3.2] and [54, §4.1] and their references therein for a thorough discussion of these operators, including their historical origin and their relation with each other.
@@
The previously mentioned Hodge theory [48, §7.4] also provides $\mathrm{W}^{2,p}$-regularity ($1<p<\infty$) for solutions of the Hodge Laplacian $\Delta_{H}$ operators (for $1$-forms) in manifolds of class $C^{2,1}$. By virtue of the musical isomorphism, we expect the result on $\Delta_{H}$ to extend to tangent vector fields, although we are unable to pinpoint a specific reference for the proof. Consequently, said regularity would also be enjoyed by the Bochner Laplacian $\Delta_{B}$, as it only differs from the Hodge Laplace $\Delta_{H}$ in lower-order terms (cf. (1.3)).
@@
If we consider the Hodge Laplacian $\Delta_{H}$ or the Surface Diffusion $\Delta_{S}$ in lieu of the Bochner Laplacian $\Delta_{B}$, different Stokes systems arise; see for instance, [36, §§4-5] for $\Delta_{S}$, and [60, §17.A] for $\Delta_{H}$.
The weak formulation for the surface diffusion operator $\Delta_{S}$ [36, eq. (4.5)] replaces the diffusion term $\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$ in (1.4) by $\int_{\Gamma}D_{\Gamma}(\mathbf{u}):D_{\Gamma}(\mathbf{v})$. Consequently, the flow velocity $\mathbf{u}$ needs to be sought in the orthogonal complement of the non-trivial kernel of the operator $D_{\Gamma}$ so as to ensure uniqueness. The well-posedness of said formulation follows from a straightforward application of the Babuška–Brezzi theory (see, e.g. [29, Th. 2.1]), which hinges on a Korn-type inequality [36, Lemma 4.1] and an inf-sup condition on the covariant divergence $\operatorname{div}_{\Gamma}$ [36, Lemma 4.2]. As expected, the well-posedness of (1.4) will depend on the same Poincaré inequality (instead of a Korn inequality) for the covariant gradient $\nabla_{\Gamma}$ in the space $\mathbf{H}^{1}_{t}(\Gamma)$ used to establish the well-posedness of (1.2).
In flat bounded domains, estimates in $\mathrm{W}^{m,p}$-norms for Stokes systems, provided such regular-enough solutions exist, are a consequence of the results by Agmon, Douglis and Nirenberg [22, 2], whose proofs are built upon techniques of integral representation and potential theory. The existence of such solutions was proved in [16] for $d=3$, while the case $d=2$ is much simpler as it reduces to a biharmonic problem [63, Chapter 1, Proposition 2.3]. We refer to [63, Chapter 1, §2.5] for a thorough discussion on the regularity of the Stokes system and its early historical development.
The special case of $\mathrm{L}^{2}$-based (i.e. $p=2$) estimates allows alternative proofs that circumvent the use of potential theory; among them we find [31, 3, 4, 32]. Ghidaglia [31] (see also [18, §3]) argues based on difference quotients and flattening of the boundary, while Amrouche and Girault [3] develop a Helmholtz decomposition approach of the space $\mathbf{W}^{m,p}(\Omega)\cap\mathbf{W}^{1,p}_{0}(\Omega)$. Finally, we refer to Giaquinta and Modica [32] for a clear and self-contained discussion of the $\mathrm{L}^{2}$-based Sobolev and Hölder regularity of nonlinear Stokes systems with general coefficients.
We highlight the particularly elegant argument by Beirão da Veiga [4] (later extended by Berselli [8]) that, via the “method of continuity” (homotopy argument) [33, §9, Theorem 5.2], transfers the $\mathbf{W}^{2,p}(\Omega)$-well-posedness of the Poisson’s equation to $\mathbf{W}^{2,p}(\Omega)\times\mathrm{W}^{1,p}(\Omega)$-well-posedness of the Stokes system.